# MODULI: Unlocking Preference Generalization via Diffusion Models for Offline Multi-Objective Reinforcement Learning

### v1

@@ 6 Conclusion
In this paper, we propose MODULI, a diffusion planning framework along with two corresponding multi-objective normalization schemes, enabling preference-return condition-guided trajectory generation for decision-making. To achieve better generalization in OOD preferences, MODULI includes a sliding guidance mechanism, training an additional slider adapter for active distribution adjustment. We conduct extensive experiments on Complete, Shattered, and Narrow types of datasets, demonstrating the superior expressive and generalization capabilities of MODULI. The results show that MODULI is an excellent Pareto front approximator, capable of expanding the Pareto front and obtaining advanced and dense solutions. But there are still limitations such as experiments being conducted only in structured state space and continuous actions. We hope to extend it to multi-objective tasks with image inputs, further enhancing its generalization ability. Additionally, nonlinear preference spaces may present new challenges.
@@ A.1 D4MORL Benchmark
In this paper, we use the Datasets for Multi-Objective Reinforcement Learning (D4MORL) [50] for experiments, a large-scale benchmark for Offline MORL. This benchmark consists of offline trajectories from 6 multi-objective MuJoCo environments, including 5 environments with 2 objectives each, namely MO-Ant, MO-HalfCheetah, MO-Hopper, MO-Swimmer, MO-Walker2d, and one environment with three objectives: MO-Hopper-3obj. The objectives in each environment are conflicting, such as high-speed running and energy saving. In D4MORL, Each environment uses PGMORL’s [47] behavioral policies for data collection. PGMORL trains a set of policies using evolutionary algorithms to approximate the Pareto front, which can map to the closest preference in the preference set for any given preference. Therefore, in D4MORL, two different quality datasets are defined: the Expert Dataset, which is collected entirely using the best preference policies in PGMORL for decision making, and the Amateur Dataset, which has a probability $P$ of using perturbed preference policies and a probability $1-p$ of being consistent with the expert dataset, with $p=0.65$. We can simply consider the Amateur Dataset as a relatively lower quality dataset with random perturbations added to the Expert Dataset. Each trajectory in the D4MORL dataset is represented as: $\tau=[\bm{\omega},\bm{s}_{0},\bm{a}_{0},\bm{r}_{0},\cdots,\bm{s}_{T},\bm{a}_{T},\bm{r}_{T}]$, where $T$ is episode length.
@@ A.3 Baselines
In this paper, we compare with various strong baselines: MODT(P) [6], MORvS(P) [12], BC(P), and MOCQL(P) [25], all of which are well-studied baselines under the PEDA [50], representing different paradigms of learning methods. All baselines were trained using the default hyperparameters and official code as PEDA. MODT(P) extends DT architectures to include preference tokens and vector-valued returns. Additionally, MODT(P) also concatenates $\bm{\omega}$ to state tokens $[\bm{s},\bm{\omega}]$ and action tokens $[\bm{a},\bm{\omega}]$ for training. MORvS(P) and MODT(P) are similar, extending the MLP structure to adapt to MORL tasks. MORvS(P) concatenates the preference with the states and the average RTGs by default as one single input. BC(P) serves as an imitation learning paradigm to train a mapping from states with preference to actions. BC(P) simply uses supervised loss and MLPs as models. MOCQL(P) represents the temporal difference learning paradigm, which learns a preference-conditioned Q-function. Then, MOCQL(P) uses underestimated Q functions for actor-critic learning, alleviating the OOD phenomenon in Offline RL.
@@ A.4 Evaluation Metrics
Definition 3 (Return Deviation (RD)).
@@
where $\bm{R}_{\psi}$ is a generalized return predictor to predict the maximum expected return achievable for a given preference $\bm{\omega}$. To accurately evaluate RD, we consistently train this network $\bm{R}_{\psi}$ using the Complete-Expert dataset, as shown in eq. 7. In this way, for any preference of a missing dataset, we can predict and know the oracle’s solution. Please note that $\bm{R}_{\psi}$ here is only used to evaluate algorithm performance and is unrelated to the Preference Predicted Normalization method in MODULI. $\bm{\omega}_{O(k)}$ is the corresponding preference of $O(k)$. A lower RD is better, as it indicates that the performance of the solution is closer to the predicted maximum under OOD preference, suggesting better generalization ability.
@@ Appendix B Implementation Details
  • We utilize a DiT structure similar to AlignDiff [10] as the backbone for all diffusion models, with an embedding dimension of 128, 6 attention heads, and 4 DiT blocks. We use a 2-layer MLP for condition embedding with inputs $[\bm{\omega},\bm{g}]$. For Preference Prediction Normalization, we use a 3-layer MLP to predict the maximum expected RTG for given preferences. For Neighborhood Preference Normalization, The neighborhood hyperparameter $\epsilon$ is fixed at $31e-$ 3 for 2-obj experiments and $33e-$ 3 for 3-obj experiments. For the slider adapter, we use a DiT network that is the same as the diffusion backbone of the slider. We observe that using the Mish activation [34] is more stable in OOD preference regions compared to the ReLU activation, which tends to diverge quickly. Finally, the model size of MODULI is around 11M.
  • All models utilize the AdamW [30] optimizer with a learning rate of $42e-$ 4 and weight decay of $51e-$ 5. We perform a total of 200K gradient updates with a batch size of 64 for all experiments. We employ the exponential moving average (EMA) model with an ema rate of $0.995$.
  • For all diffusion models, we use a continuous linear noise schedule [37] for $\alpha_{s}$ and $\sigma_{s}$. We employ DDIM [41] for denoising generation and use a sampling step of $10$. For all tasks, we use a sampling temperature of $0.5$.
@@
  • The inverse dynamic model is implemented as a 3-layer MLP. The first two layers consist of a Linear layer followed by a Mish [34] activation and a LayerNorm. And, the final layer is followed by a Tanh activation. This model utilizes the AdamW optimizer and is trained for 500K gradient steps.
@@ Appendix C In-depth Analysis of Different Normalization Methods
As shown in fig. 6, we visualize the trajectory returns of the Walker2d-amateur-Complete dataset after applying different normalization methods. In order to achieve refined guided generation, we require consistency between the training objectives and the sampling objectives, meaning the sampling objectives should correspond to high-quality trajectories in the dataset. A simple Global Normalization fails to achieve this, as it cannot identify a definitive high-return guiding objective that matches the return distribution of dataset. Setting the return to $\bm{1}^{\bm{n}}$ may be unachievable for Global Normalization. On the contrary, Preference Predicted Normalization (PPN) and Neighborhood Preference Normalization (NPN) successfully normalize the trajectories in the dataset into distinctly different trajectories. High-quality trajectories that conform to preferences are clustered around $\bm{1}^{\bm{n}}$, and trajectories points farther from $\bm{1}^{\bm{n}}$ represent lower vector returns. Note that at this time, the training objective and the sampling objective are fully aligned. Compared with NPN, the PPN method uses neural networks to predict the maximum expected RTG for each preference, facing prediction errors in low-quality datasets (e.g. amateur dataset). After PPN process using $\bm{g}_{\tau}/\bm{g}_{\text{max}}(\bm{\omega})$, a small number of trajectories even slightly exceed $\bm{1}^{\bm{n}}$, indicating that $\bm{R}_{\psi}(\bm{\omega})$ fails to accurately predict the maximum expected return. This ultimately causes the guidance target $\bm{1}^{\bm{n}}$ used in the deployment phase to mismatch the target corresponding to the high trajectory return in training, resulting in a suboptimal policy. The NPN method can achieve accurate normalization for multi-objective targets, significantly distinguishing the quality of trajectories in the dataset, while ensuring that the sampling targets during the training phase and the deployment phase are consistent.
@@ Appendix D Proofs for Sliding Guidance
$p_{\theta}\left(\bm{x}^{0}(\tau)\mid c\right)\leftarrow p_{\theta}\left(\bm{x}^{0}(\tau)\mid c\right)\left(\frac{p_{\theta^{*}}\left(c_{+}\mid\bm{x}^{0}(\tau)\right)}{p_{\theta^{*}}\left(c_{-}\mid\bm{x}^{0}(\tau)\right)}\right)^{\eta}$ (14)
where $p_{\theta}\left(\bm{x}^{0}(\tau)\mid c\right)$ represents the distribution generated by the original diffusion model $\theta$ when conditioned on $c$. According to the Bayesian formula $p\left(c_{+}\mid\bm{x}^{0}(\tau)\right)=\frac{p\left(\bm{x}^{0}(\tau)\mid c_{+}\right)p\left(c_{+}\right)}{p\left(\bm{x}^{0}(\tau)\right)}$, the gradient of the log probability $\nabla\log{[p_{\theta}\left(\bm{x}^{0}(\tau)\mid c\right)\left(\frac{p_{\theta^{*}}\left(c_{+}\mid\bm{x}^{0}(\tau)\right)}{p_{\theta^{*}}\left(c_{-}\mid\bm{x}^{0}(\tau)\right)}\right)^{\eta}]}$ can be expanded to:
$\nabla\log{[p_{\theta}\left(\bm{x}^{0}(\tau)\mid c\right)\left(\frac{p_{\theta^{*}}\left(\bm{x}^{0}(\tau)\mid c_{+}\right)p\left(c_{+}\right)}{p_{\theta^{*}}\left(\bm{x}^{0}(\tau)\mid c_{-}\right)p\left(c_{-}\right)}\right)^{\eta}]}$ (15)
@@
$\nabla\log{p_{\theta}\left(\bm{x}^{0}(\tau)\mid c\right)}+\eta\left(\nabla\log{p_{\theta^{*}}\left(\bm{x}^{0}(\tau)\mid c_{+}\right)}-\nabla\log{p_{\theta^{*}}\left(\bm{x}^{0}(\tau)\mid c_{-}\right)}\right)$ (16)
Based on the reparametrization trick of [19], we can introduce a time-varying noising process and express each score (gradient of log probability) as a denoising prediction $\epsilon\left(\bm{x}^{t},c,t\right)$, thus eq. 16 is equivalent to modifying the noise prediction model:
@@ Appendix E Pseudo Code
The training process of the MODULI consists of two parts: the training of the diffusion model and the training of the slider adapter, which are described in Algorithms 1 and 2, respectively. Given a normalized dataset $D_{g}$ containing trajectories of states $\bm{s}$, normalized RTG $\bm{g}$ and preferences $\bm{\omega}$, MODULI first trains a diffusion planner $\theta$:
@@
Appendix F Full Results
@@
Appendix G Additional Experiment Results
G.1 Performance with Different Guidance Mechanism
@@
G.2 Approximate Pareto Front on all Environments
We present the approximate Pareto front visualization results of MODULI for all environments in fig. 7. MODULI is a good Pareto front approximator, capable of extending the front on the amateur dataset and performing interpolation or extrapolation generalization on the Shattered or Narrow datasets.

### v2

@@ 6 Conclusion
In this paper, we propose MODULI, a diffusion planning framework with two multi-objective normalization schemes, enabling preference-return condition-guided trajectory generation for decision-making. To achieve better generalization in OOD preferences, MODULI includes a sliding guidance mechanism, training an additional slider adapter for active distribution adjustment. We conduct extensive experiments on Complete, Shattered, and Narrow datasets, demonstrating the superior expressive and generalization capabilities of MODULI. The results show that MODULI is an excellent Pareto front approximator, capable of expanding the Pareto front and obtaining advanced, dense solutions. However, there are limitations, such as experiments conducted in low-dimensional state space and continuous action space. We hope to extend it to multi-objective tasks with image inputs, further enhancing its generalization ability. Additionally, nonlinear preference spaces may present new challenges.
Acknowledge
This work is supported by the National Natural Science Foundation of China (Grant Nos. 62422605, 92370132). We would like to thank all the anonymous reviewers for their valuable comments and constructive suggestions, which have greatly improved the quality of this paper.
@@ A.1 D4MORL Benchmark
In this paper, we use the Datasets for Multi-Objective Reinforcement Learning (D4MORL) (Zhu et al., 2023a) for experiments, a large-scale benchmark for Offline MORL. This benchmark consists of offline trajectories from 6 multi-objective MuJoCo environments, including 5 environments with 2 objectives each, namely MO-Ant, MO-HalfCheetah, MO-Hopper, MO-Swimmer, MO-Walker2d, and one environment with three objectives: MO-Hopper-3obj. The objectives in each environment are conflicting, such as high-speed running and energy saving. In D4MORL, Each environment uses PGMORL’s (Xu et al., 2020) behavioral policies for data collection. PGMORL trains a set of policies using evolutionary algorithms to approximate the Pareto front, which can map to the closest preference in the preference set for any given preference. Therefore, in D4MORL, two different quality datasets are defined: the Expert Dataset, which is collected entirely using the best preference policies in PGMORL for decision making, and the Amateur Dataset, which has a probability $P$ of using perturbed preference policies and a probability $1-p$ of being consistent with the expert dataset, with $p=0.65$. We can simply consider the Amateur Dataset as a relatively lower quality dataset with random perturbations added to the Expert Dataset. Each trajectory in the D4MORL dataset is represented as: $\tau=[\bm{\omega},\bm{s}_{0},\bm{a}_{0},\bm{r}_{0},\cdots,\bm{s}_{T},\bm{a}_{T},\bm{r}_{T}]$, where $T$ is episode length.
@@ A.3 Baselines
In this paper, we compare with various strong baselines: MODT(P) (Chen et al., 2021), MORvS(P) (Emmons et al., 2021), BC(P), and MOCQL(P) (Kumar et al., 2020), all of which are well-studied baselines under the PEDA (Zhu et al., 2023a), representing different paradigms of learning methods. All baselines were trained using the default hyperparameters and official code as PEDA. MODT(P) extends DT architectures to include preference tokens and vector-valued returns. Additionally, MODT(P) also concatenates $\bm{\omega}$ to state tokens $[\bm{s},\bm{\omega}]$ and action tokens $[\bm{a},\bm{\omega}]$ for training. MORvS(P) and MODT(P) are similar, extending the MLP structure to adapt to MORL tasks. MORvS(P) concatenates the preference with the states and the average RTGs by default as one single input. BC(P) serves as an imitation learning paradigm to train a mapping from states with preference to actions. BC(P) simply uses supervised loss and MLPs as models. MOCQL(P) represents the temporal difference learning paradigm, which learns a preference-conditioned Q-function. Then, MOCQL(P) uses underestimated Q functions for actor-critic learning, alleviating the OOD phenomenon in Offline RL.
@@ A.4 Evaluation Metrics
Definition A.3 (Return Deviation (RD)).
@@
where $\bm{R}_{\psi}$ is a generalized return predictor to predict the maximum expected return achievable for a given preference $\bm{\omega}$. To accurately evaluate RD, we consistently train this network $\bm{R}_{\psi}$ using the Complete-Expert dataset, as shown in Equation 9. In this way, for any preference of a missing dataset, we can predict and know the oracle’s solution. Please note that $\bm{R}_{\psi}$ here is only used to evaluate algorithm performance and is unrelated to the Preference Predicted Normalization method in MODULI. $\bm{\omega}_{O(k)}$ is the corresponding preference of $O(k)$. A lower RD is better, as it indicates that the performance of the solution is closer to the predicted maximum under OOD preference, suggesting better generalization ability.
@@ Appendix B Implementation Details
  • We utilize a DiT structure similar to AlignDiff (Dong et al., 2023) as the backbone for all diffusion models, with an embedding dimension of 128, 6 attention heads, and 4 DiT blocks. We use a 2-layer MLP for condition embedding with inputs $[\bm{\omega},\bm{g}]$. For Preference Prediction Normalization, we use a 3-layer MLP to predict the maximum expected RTG for given preferences. For Neighborhood Preference Normalization, The neighborhood hyperparameter $\epsilon$ is fixed at $31e-$ 3 for 2-obj experiments and $33e-$ 3 for 3-obj experiments. For the slider adapter, we use a DiT network that is the same as the diffusion backbone of the slider. We observe that using the Mish activation (Misra, 2019) is more stable in OOD preference regions compared to the ReLU activation, which tends to diverge quickly. Finally, the model size of MODULI is around 11M.
  • All models utilize the AdamW (Loshchilov & Hutter, 2017) optimizer with a learning rate of $42e-$ 4 and weight decay of $51e-$ 5. We perform a total of 200K gradient updates with a batch size of 64 for all experiments. We employ the exponential moving average (EMA) model with an ema rate of $0.995$.
  • For all diffusion models, we use a continuous linear noise schedule (Nichol & Dhariwal, 2021) for $\alpha_{s}$ and $\sigma_{s}$. We employ DDIM (Song et al., 2021a) for denoising generation and use a sampling step of $10$. For all tasks, we use a sampling temperature of $0.5$.
@@
  • The inverse dynamic model is implemented as a 3-layer MLP. The first two layers consist of a Linear layer followed by a Mish (Misra, 2019) activation and a LayerNorm. And, the final layer is followed by a Tanh activation. This model utilizes the AdamW optimizer and is trained for 500K gradient steps.
@@
  • We standardized the parameter size of all baseline algorithms to the same magnitude as much as possible. MORvS: 8.55M, MOBC:8.83M, MODT:9.75M.
@@ Appendix C In-depth Analysis of Different Normalization Methods
As shown in Figure 6, we visualize the trajectory returns of the Walker2d-amateur-Complete dataset after applying different normalization methods. In order to achieve refined guided generation, we require consistency between the training objectives and the sampling objectives, meaning the sampling objectives should correspond to high-quality trajectories in the dataset. A simple Global Normalization fails to achieve this, as it cannot identify a definitive high-return guiding objective that matches the return distribution of dataset. Setting the return to $\bm{1}^{\bm{n}}$ may be unachievable for Global Normalization. On the contrary, Preference Predicted Normalization (PPN) and Neighborhood Preference Normalization (NPN) successfully normalize the trajectories in the dataset into distinctly different trajectories. High-quality trajectories that conform to preferences are clustered around $\bm{1}^{\bm{n}}$, and trajectories points farther from $\bm{1}^{\bm{n}}$ represent lower vector returns. Note that at this time, the training objective and the sampling objective are fully aligned. Compared with NPN, the PPN method uses neural networks to predict the maximum expected RTG for each preference, facing prediction errors in low-quality datasets (e.g. amateur dataset). After PPN process using $\bm{g}_{\tau}/\bm{g}_{\text{max}}(\bm{\omega})$, a small number of trajectories even slightly exceed $\bm{1}^{\bm{n}}$, indicating that $\bm{R}_{\psi}(\bm{\omega})$ fails to accurately predict the maximum expected return. This ultimately causes the guidance target $\bm{1}^{\bm{n}}$ used in the deployment phase to mismatch the target corresponding to the high trajectory return in training, resulting in a suboptimal policy. The NPN method can achieve accurate normalization for multi-objective targets, significantly distinguishing the quality of trajectories in the dataset, while ensuring that the sampling targets during the training phase and the deployment phase are consistent.
@@ Appendix D Proofs for Sliding Guidance
$p_{\theta}\left(\boldsymbol{x}^{0}(\tau)\mid c\right)\leftarrow p_{\theta}\left(\boldsymbol{x}^{0}(\tau)\mid c\right)\left(\frac{p_{\theta^{*}}\left(c_{+}\mid\boldsymbol{x}^{0}(\tau)\right)}{p_{\theta^{*}}\left(c_{-}\mid\boldsymbol{x}^{0}(\tau)\right)}\right)^{\eta}$ (1)
where $p_{\theta}\left(\boldsymbol{x}^{0}(\tau)\mid c\right)$ represents the distribution generated by the original diffusion model $\theta$ when conditioned on $c$. According to the Bayesian formula $p\left(c_{+}\mid\boldsymbol{x}^{0}(\tau)\right)=\frac{p\left(\boldsymbol{x}^{0}(\tau)\mid c_{+}\right)p\left(c_{+}\right)}{p\left(\boldsymbol{x}^{0}(\tau)\right)}$, the gradient of the log probability $\nabla\log{[p_{\theta}\left(\boldsymbol{x}^{0}(\tau)\mid c\right)\left(\frac{p_{\theta^{*}}\left(c_{+}\mid\boldsymbol{x}^{0}(\tau)\right)}{p_{\theta^{*}}\left(c_{-}\mid\boldsymbol{x}^{0}(\tau)\right)}\right)^{\eta}]}$ can be expanded to:
$\nabla\log{[p_{\theta}\left(\boldsymbol{x}^{0}(\tau)\mid c\right)\left(\frac{p_{\theta^{*}}\left(\boldsymbol{x}^{0}(\tau)\mid c_{+}\right)p\left(c_{+}\right)}{p_{\theta^{*}}\left(\boldsymbol{x}^{0}(\tau)\mid c_{-}\right)p\left(c_{-}\right)}\right)^{\eta}]}$ (2)
@@
$\nabla\log{p_{\theta}\left(\boldsymbol{x}^{0}(\tau)\mid c\right)}+\eta\left(\nabla\log{p_{\theta^{*}}\left(\boldsymbol{x}^{0}(\tau)\mid c_{+}\right)}-\nabla\log{p_{\theta^{*}}\left(\boldsymbol{x}^{0}(\tau)\mid c_{-}\right)}\right)$ (3)
Based on the reparametrization trick of (Ho et al., 2020), we can introduce a time-varying noising process and express each score (gradient of log probability) as a denoising prediction $\epsilon\left(\bm{x}^{t},c,t\right)$, thus Equation 3 is equivalent to modifying the noise prediction model:
@@ Appendix E Pseudo Code
The training process of the MODULI consists of two parts: the training of the diffusion model and the training of the slider adapter, which are described in Algorithm 1 and Algorithm 2, respectively. Given a normalized dataset $D_{g}$ containing trajectories of states $\bm{s}$, normalized RTG $\bm{g}$ and preferences $\bm{\omega}$, MODULI first trains a diffusion planner $\theta$:
@@
Appendix F Details of Loss-weight Trick
The diffusion planner generates a state sequence $x^{0}(\tau)=[s_{0},\cdots,s_{H-1}]$. Since closed-loop control is used, only $s_{0}$ and $s_{1}$ are most relevant to the current decision. When calculating the loss, the weight of $s_{1}$ is increased from 1 to 10, which is considered more important. For clarity, we provide pseudocode below for the loss calculation of the diffusion model incorporating the loss-weight trick and explain it in detail.
# Create a loss_weight and assign a higher
loss weight to the next observation (o_1).
next_obs_loss_weight = 10
loss_weight = torch.ones((batch_size, horizon, obs_dim))
loss_weight[:, 1, :] = next_obs_loss_weight
def diffusion_loss_pseudocode(x0, condition):
Pseudocode function to demonstrate the loss function in diffusion models
Variable descriptions:
x0: Original noise-free data - [batch_size, horizon, obs_dim]
condition: Conditional information - [batch_size, condition_dim]
xt: Noisy data after noise addition - same shape as x0
t: Time points in the diffusion process step - [batch_size]
eps: Noise added to x0 - same shape as x0
loss_weight: Loss weight coefficient - scalar or tensor matching x0’s shape
# Step 1: Add noise to the original data
xt, t, eps = add_noise(x0) # Generate noisy xt from x0
# Step 2: Process conditional information
# Encode the original condition information
condition = condition_encoder(condition)
# Step 3: Predict noise through the diffusion model
predicted_eps = diffusion_model(xt, t, condition) # Model predicts the added noise
# Step 4: Calculate mean
squared error
# Mean squared error between predicted noise and actual noise
loss = (predicted_eps - eps) ** 2
# Step 5: Apply loss weights
# We assign a higher weight to the prediction of the s_1
loss = loss * loss_weight
# Step 6: Calculate average loss
final_loss = loss.mean() # Calculate the average loss
return final_loss
Appendix G Full Results
@@
Appendix H Additional Experiment Results
H.1 Performance with Different Guidance Mechanism
@@
H.2 Approximate Pareto Front on all Environments
We present the approximate Pareto front visualization results of MODULI for all environments in Figure 7. MODULI is a good Pareto front approximator, capable of extending the front on the amateur dataset and performing interpolation or extrapolation generalization on the Shattered or Narrow datasets.